# AANet: Virtual Screening under Structural Uncertainty via Alignment and Aggregation

### v1

@@ B.6 Docking on LIT-PCBA
B.7 Baseline implementation details
@@
RTMScore [21]: Implementation obtained from the official GitHub https://github.com/sc8668/RTMScore and their pretrained weights on the PDBbind-v2020 [30] general set. The authors excluded any complexes overlapping with the PDBbind-v2020 core set and the CASF-2016 [39] benchmark (reducing from 19,443 to 19,149 complexes), but did not apply filtering against DUD-E or LIT-PCBA. Following the repository defaults, pockets are defined as all receptor atoms within 10 Å of the ligand, or the cavity in the annotated setting, along with the default scoring scheme.
EquiScore [20]: We downloaded the code and model weights (trained on PDBscreen, with all proteins from DUD-E and DEKOIS2.0 [40] excluded from the training data, but not those from LIT-PCBA) from https://github.com/CAODH/EquiScore. Each molecule’s own docking pose was used to define its pocket, i.e., receptor residues within 8 Å. Scores were then obtained using the default inference configuration.
@@
TankBind [22]: We downloaded the official implementation from https://github.com/luwei0917/TankBind along with their pretrained weights on PDBbind-v2020 [30]. Following the EquiBind [41] time split, complexes deposited before 2019 (17,787 after RDKit filtering) constituted the training set, those from 2019 (968) the validation set, and those after 2019 (363) the test set, without deduplication against DUD-E or LIT-PCBA. We followed the original settings for all preprocessing and hyperparameters. Pockets were detected using P2Rank [27], with a 20 Å radius around each predicted center. In the annotated setting, when a true pocket was available, we computed its centroid and selected the single P2Rank pocket whose center was closest. In the blind setting, where no ligand was provided, we scored all P2Rank pockets and retained the maximum predicted affinity per compound.
DrugCLIP [3]: We used the official DrugCLIP implementation from https://github.com/bowen-gao/DrugCLIP. The model was fine-tuned on the PDBbind-v2019 general set, with all complexes overlapped with those in DUD-E or LIT-PCBA removed to ensure zero-shot evaluation. All other hyperparameters and data processing followed the authors’ original settings.
@@ Appendix C Additional results
C.1 AUROC under annotated and blind settings
@@
Table S7: AUROC performance on DUD-E and LIT-PCBA. Each method is evaluated on three structural subsets: holo, apo (experimental), and apo (predicted), under both annotated and blind settings. Row colors indicate method type: Docking & Rescoring, Docking-free baseline, and Proposed method.
Method Holo apo-exp (annot) apo-pred (annot) apo-exp (blind) apo-pred (blind)
@@
Glide-SP 0.6119 0.5533 0.5725
RTMScore 0.7211 0.6077 0.6246
EquiScore 0.7257 0.6478 0.6631
TankBind 0.7872 0.7765 0.8006 0.7745 0.7897
DrugCLIP 0.8133 0.7518 0.7826 0.6134 0.5798
AANet 0.8963 0.8744 0.8752 0.8593 0.8617
@@
Glide-SP 0.5078 0.5070 0.5037
RTMScore 0.5245 0.5161 0.5324
EquiScore 0.5800 0.4694 0.5342
TankBind 0.6110 0.5816 0.6121 0.5920 0.5942
DrugCLIP 0.5742 0.5851 0.5635 0.5006 0.4805
AANet 0.5621 0.5737 0.5640 0.5750 0.5641
C.2 Results for holo full sets
@@
AUROC BEDROC ($\alpha=80.5$) EF 1% AUROC BEDROC ($\alpha=80.5$) EF 1%

### v2

@@ B.7 Docking on LIT-PCBA
B.8 Baseline implementation details
@@
RTMScore [22]: Implementation obtained from the official GitHub https://github.com/sc8668/RTMScore and their pretrained weights on the PDBbind‑v2020 [31] general set. The authors excluded any complexes overlapping with the PDBbind‑v2020 core set and the CASF‑2016 [40] benchmark (reducing from 19,443 to 19,149 complexes), but did not apply filtering against DUD-E or LIT‑PCBA. Following the repository defaults, pockets are defined as all receptor atoms within 10 Å of the ligand, or the cavity in the annotated setting, along with the default scoring scheme.
EquiScore [21]: We downloaded the code and model weights (trained on PDBscreen, with all proteins from DUD‑E and DEKOIS2.0 [41] excluded from the training data, but not those from LIT‑PCBA) from https://github.com/CAODH/EquiScore. Each molecule’s own docking pose was used to define its pocket, i.e., receptor residues within 8 Å. Scores were then obtained using the default inference configuration.
@@
TankBind [23]: We downloaded the official implementation from https://github.com/luwei0917/TankBind along with their pretrained weights on PDBbind‑v2020 [31]. Following the EquiBind [42] time split, complexes deposited before 2019 (17,787 after RDKit filtering) constituted the training set, those from 2019 (968) the validation set, and those after 2019 (363) the test set, without deduplication against DUD-E or LIT‑PCBA. We followed the original settings for all preprocessing and hyperparameters. Pockets were detected using P2Rank [28], with a 20 Å radius around each predicted center. In the annotated setting, when a true pocket was available, we computed its centroid and selected the single P2Rank pocket whose center was closest. In the blind setting, where no ligand was provided, we scored all P2Rank pockets and retained the maximum predicted affinity per compound.
DrugCLIP [3]: We used the official DrugCLIP implementation from https://github.com/bowen-gao/DrugCLIP. The model was fine‐tuned on the PDBbind-v2019 general set, with all complexes overlapped with those in DUD‑E or LIT‑PCBA removed to ensure zero‐shot evaluation. All other hyperparameters and data processing followed the authors’ original settings.
@@ Appendix C Additional results
C.1 Benchmarking computational cost on the LIT-PCBA dataset
As our dual-tower backbone does not involve fusion prior to the adapter and supports embedding reuse, AANet introduces minimal overhead beyond DrugCLIP while providing significantly improved performance, particularly in unbound settings. As shown in Table S7, the majority of runtime during the scoring phase arises from data loading and CPU–GPU communication, rather than model inference itself. On typical commercial-scale libraries, AANet maintains ultra-fast screening throughput comparable to DrugCLIP.
Table S7: Comparison of computational cost across methods. *Sharing cost refers to preprocessing that can be reused. **PCBA-One corresponds to the subset of 2.8M complexes.
Method Sharing Cost* PCBA-One** (2.8M) PCBA-Full (10M) PCBA-Apo-Blind (100M) Overall Cost
Glide LigPrep: 1 hr (930 mols/s) 3 days 11 days 3.6 months 3 days – 3.6 months
RTMScore
Preprocessing: 3 hrs (247/s)
Inference: 8.2 hrs (94/s)
11.2 hrs + 29.5 hrs 17 days 11.2 hrs – 17 days
EquiScore
Preprocessing: 2.7 hrs (291/s)
Inference: 8.1 hrs (96/s)
11 hrs + 29 hrs 16.7 days 10.8 hrs – 16.7 days
TankBind 7 hrs 25 hrs 10.4 days 7 hrs – 10.4 days
DrugCLIP
Mol: 270 s (10,370/s)
Pocket: 1–2 s (86/s)
15 s 15 s $\sim$300 s
AANet
Mol: 270 s (10,370/s)
Pocket: 1–3 s (36/s)
15 s 15 s $\sim$300 s
C.2 AUROC under annotated and blind settings
@@
Table S8: AUROC performance on DUD‑E and LIT‑PCBA. Each method is evaluated on three structural subsets: holo, apo (experimental), and apo (predicted), under both annotated and blind settings. Row colors indicate method type: Docking & Rescoring, Docking-free baseline, and Proposed method.
Method Holo
apo-exp
apo-pred
apo-exp
apo-pred
@@
[…]
@@
[…]
C.3 Results for holo full sets
@@
AUROC
BEDROC
($\alpha=80.5$)
EF 1% AUROC
BEDROC
($\alpha=80.5$)
EF 1%
@@
C.4 Results on the DEKOIS dataset
we mapped the DEKOIS targets to UniProt ID and successfully downloaded 77 target structures from the AlphaFold Database, excluding 4 virus-related targets that are not available (HIV1PR & HIV1RT: P03366, SARS-HCoV: P0C6X7, NA: P27907).
As shown in Table S10, our method consistently outperforms DrugCLIP across all settings, and surpasses traditional docking baselines under holo pocket settings (italic) when tested under the most challenging apo-pred blind setup (bolded).
Table S10: Performance on DEKOIS (77 targets) under holo, apo-pred-annotated (the cavity detected by FPocket that is closest to the holo pocket), and apo-pred-blind settings.
Method
AUROC
(holo) $\uparrow$
AUROC
(apo-annot)
AUROC
(apo-blind)
BEDROC
(holo) $\uparrow$
BEDROC
(apo-annot)
BEDROC
(apo-blind)
EF@0.5%
(holo) $\uparrow$
EF@0.5%
(apo-annot)
EF@0.5%
(apo-blind)
Vina 2.58
GOLD 10.32
Glide 11.85
DrugCLIP 0.7967 0.7075 0.4355 0.532 0.322 0.0313 20.62 12.12 0.92
AANet 0.8785 0.8624 0.7708 0.6446 0.5824 0.3891 23.26 20.91 13.33
C.5 Results on the DUD-E dataset under de-redundant / de-homologous setting
On PDBbind, we removed duplicate complexes based on structure IDs, following prior work to ensure fair comparison. However, this may not fully eliminate target-level redundancy. For some of baselines, duplicates were not removed (Appendex B.8). On ChEMBL, we filtered data using UniProt IDs, which more reliably removes overlapping targets.
To further address potential leakage, we conducted an additional experiment by using MMseqs2 [43] to identify and remove any training proteins (from PDBbind and ChEMBL) containing chains with $\geq$ 90% sequence identity to any DUD-E structure / target sequences. We then compared our method and DrugCLIP under both the original and this more stringent de-redundant / de-homologous setting. As shown in Table S11, our method maintains robust performance even after removing homologous sequences, demonstrating better generalization.
Table S11: Performance on DUD-E under de-redundant and de-homologous (90%) settings.
Method
BEDROC
BEDROC
(apo-pred annot)
BEDROC
(apo-exp annot)
BEDROC
(apo-pred blind)
BEDROC
(apo-exp blind)
EF@1%
EF@1%
(apo-pred annot)
EF@1%
(apo-exp annot)
EF@1%
(apo-pred blind)
EF@1%
(apo-exp blind)
DrugCLIP
(de-redundant)
0.5157 0.3746 0.3493 0.1974 0.1926 33.70 22.70 21.36 12.05 11.75
DrugCLIP
(de-homologous)
0.3949 0.2323 0.2134 0.1097 0.1298 24.90 13.56 12.59 6.51 7.36
AANet
(de-redundant)
0.6365 0.6003 0.5866 0.6232 0.5764 40.85 38.46 38.03 40.85 37.19
AANet
(de-homologous)
0.5985 0.5247 0.5475 0.5955 0.5411 38.55 33.59 34.76 38.34 34.50
C.6 Comparison with ligand-based methods
ligand-based virtual screening (LBVS) can bypass the need for protein structural information and is particularly useful when high-quality ligands are available or when protein structures are unreliable.
However, LBVS is often more straightforward when good query ligands are available (e.g., potent, exogenous ligands), but may be less effective in identifying novel chemotypes or in the absence of relevant prior ligands. In contrast, SBVS leverages the protein’s structural context, which can help discover structurally diverse actives but depends on structure quality.
We evaluated several widely-used LBVS methods, covering both 2D and 3D paradigms, including ECFP (2D topology-based), USR (3D shape-based), and Shape from Schrödinger suite (3D shape overlap). The results are summarized in Table S12.
Table S12: Performance of common LBVS methods on DUD-E benchmark.
Method
EF@1%
BEDROC ($\alpha=80.5$)
USR [36] 2.74 0.051
PhaseShape [37] 8.29 0.138
ECFP4 [44] 26.61 0.414
AANet (apo-blind) 37.19 0.576
The notably strong performance of ECFP4 is likely due to the fact that many actives in virtual screening benchmarks such as DUD-E share similar topological scaffolds, whereas decoys are specifically designed to have similar physicochemical properties but divergent topologies. As a result, ECFP4 can effectively distinguish actives from decoys based on topological similarity alone, leading to higher enrichment factors. However, such advantages may not always generalize to real-world screening scenarios, where novel actives are often needed.
Appendix D Additional discussion
D.1 Selection of the pocket extraction radius
This choice of 10 Å pocket extraction radius was based on quantitative analysis on the PDBbind dataset, supported by the following three key findings:
1. Coverage of original holo pockets: The oracle 10 Å cavities, which are defined as those with the highest overlap (IoU) with holo pockets, fully cover 75.38% of the original 6 Å holo pockets, and 90.94% achieve $\geq$ 85% coverage. This indicates that 10 Å cavities provide a reliable approximation of the ground-truth pocket regions.
2. Tradeoff between feature integrity and computational cost: The average atom count in 10 Å oracle cavities is 612.68 ± 240.75, over 3× larger than the 6 Å holo pockets (199.32 ± 55.04). Since computational cost scales non-linearly with the number of input atoms, we uniformly downsample pockets to $\leq$ 256 atoms during training to ensure fair comparisons across different pocket definitions. Larger regions would require more aggressive downsampling, which may compromise spatial features or increase computational overhead. Thus, 10 Å represents a practical upper bound that balances representational fidelity with efficiency.
3. Avoiding bias from over-extension: Extending the radius beyond 10 Å risks incorporating off-target cavities, especially on cluttered protein surfaces. As our 10 Å is defined by minimal atom-to-atom distance (not center-based), it already captures ample structural context. Further expansion would dilute the binding signal and potentially introduce irrelevant patterns.
Thus, the 10 Å threshold strikes a practical balance: it ensures high coverage of the ground-truth binding pocket, preserves structural integrity, and avoids unnecessary computational overhead or learning from irrelevant patterns.
D.2 The role of holo pocket in tri-modal contrastive learning
The alignment with holo pockets plays a critical and bidirectional role in our tri-modal contrastive framework, motivated by the following:
1. Holo $\rightarrow$ Cavity - Guidance for learning from noisy input: Holo pockets, shaped by actual ligand binding, serve as reliable supervisory anchors (as stated by the reviewer). Aligning detected cavities with holo pockets helps the model extract meaningful features even from noisy or imprecise cavity detection.
2. Cavity $\rightarrow$ Holo - Addressing structural uncertainty and generalization: Holo pockets are biased toward a specific ligand series (Appendix A), while other actives may bind elsewhere. By aligning cavities (which are ligand-agnostic) to holo pockets, the model learns features that are spatially intrinsic to the protein. This improves the model’s robustness to structural uncertainty and enhances generalization. This alignment also boosts performance on holo structures, since the learned features are less tied to ligand-specific conformations.
3. Cavity $\leftrightarrow$ Holo - Dual modeling of general and specific signals: The model jointly learns ligand-independent (intrinsic to proteins) and ligand-specific pocket representations, which improves robustness. This enables stronger performance across both holo and apo structures, not just mitigating the performance gap but outperforming DrugCLIP in both cases.
In summary, holo-cavity alignment is essential for modeling both precise and generalizable binding patterns, supporting our goal of robust virtual screening under structural uncertainty.
D.3 Strengths and limitations of pocket detection
While our current framework employs Fpocket for cavity detection during training, AANet demonstrates robustness across different cavity detectors. Although our model uses Fpocket to extract training cavities, the downstream performance remains stable across different cavity detection tools (Table 4). These conventional detectors are grounded in physicochemical definitions of binding pockets, such as surface concavity, void volume, and geometric constraints. The consistency across tools suggests that AANet is not overfitting to a specific detection algorithm, but instead learns underlying structural priors intrinsic to proteins, such as spatial enclosures or clefts.
Alignment with practical structure-based pipelines: Our method follows the typical structure-based virtual screening pipeline, where pocket detection and binding assessment are distinct stages. This modularity aligns with standard workflows in drug discovery and allows our method to plug into existing infrastructure while remaining adaptable to alternative cavity detectors.
Physically meaningful supervision: By grounding learning on cavities defined via physical heuristics (e.g., void detection), we introduce inductive bias from medicinal chemistry intuition into the model. This helps prevent overfitting to spurious binding patterns and encourages the model to focus on biophysically plausible pockets. Concerns about physical plausibility have recently gained attention in the structure prediction and docking communities, as highlighted by works such as PoseBusters in deep learning-based docking. This emphasis on biophysical realism is equally relevant to virtual screening and motivates the design of our method.
Although we currently do not explore end-to-end cavity prediction, which is a promising direction, especially on novel targets including flexible proteins. As datasets grow and modeling techniques evolve, we believe this integration could offer even more robust and generalizable solutions.
Limitation in flexible or disordered proteins: We acknowledge that for highly flexible or intrinsically disordered proteins, where well-formed pockets are absent and ligand-induced fit is essential, existing pocket-based SBVS methods, including ours, face inherent challenges. Our work primarily addresses uncertainty in binding site localization, not full flexible binding modeling.